### v3

# A new sociology of humans and machines

Abstract
From fake social media accounts and generative-AI chatbots to trading algorithms and self-driving vehicles, robots, bots, and algorithms are proliferating and permeating our communication channels, social interactions, economic transactions, and transportation arteries. Networks of multiple interdependent and interacting humans and intelligent machines constitute complex social systems where the collective outcomes cannot be deduced from either human or machine behavior alone. Under this paradigm, we review recent research and identify general dynamics and patterns in situations of competition, coordination, cooperation, contagion, and collective decision-making, with context-rich examples from high-frequency trading markets, a social media platform, an open-collaboration community, and a discussion forum. To ensure more robust and resilient human-machine communities, we require a new sociology of humans and machines. Researchers should study these communities using complex-system methods, engineers should explicitly design AI for human-machine and machine-machine interactions, and regulators should govern the ecological diversity and social co-development of humans and machines.
Robotic trains and cars drive us around, auction bots outbid us for purchases, ChatGPT answers our questions, while social media bots feed us with dubious facts and news. Modern society is a complex human-machine social system in which machines are becoming more numerous, human interactions with machines – more frequent, and machine-machine interactions – more consequential. With recent advances in generative AI models, the existential threat of unexplainable and uncontrollable general AI is looming large again 1, 2, 3. However, when they are numerous and interdependent, even simple unintelligent artificial agents can produce unintended and potentially undesirable outcomes. If we want to prevent financial crashes, improve road safety, preserve market competition, increase auction market efficiency, and reduce misinformation, it is no longer sufficient to understand humans – we need to consider machines, understand how humans and machines interact, and how the collective behavior of systems of humans and machines can be predicted. We require a new sociology of humans and machines.
This Perspective synthesizes research and ideas related to social systems composed of multiple autonomous yet interacting and interdependent humans and machines such as algorithms, bots, and robots (Fig. 1). Similar to the conceptualizations of socio-technical systems 4, actor-network theory 5, 6, cyber-physical social systems 7, 8, social machines 9, 10, 11, and human-machine networks 12, 13, our principal assumption is that humans and machines form a single social system. In contrast, we do not approach machines as a single medium or entity – “technology” – but emphasize their multiplicity, independence, and heterogeneity and their interactions 14. Further, we approach the aggregates as complex social systems where network effects and nonlinear dynamic processes produce collective outcomes that cannot be necessarily deduced from individual preferences and behavior alone15. Our conceptualization extends and complements the ecological approach to studying machine behavior 16, the “hybrid collective intelligence perspective” 17, and the budding field of Social AI 18. We aim to offer a conceptual overview of the topic. Hence, we do not utilize a systematic literature search strategy and instead present selected examples to showcase the new conceptualization 19, 20, 21.
Human-machine interactions
We use the term “machines” to refer to a range of computational artifacts: embodied in physical devices such as humanoid robots or existing only in digital space, such as bots and algorithms, and varying in sophistication from simple expert systems that use pre-defined if-else rules to generative deep-learning models that learn from data in real-time. Like humans, the machines we consider exhibit diverse goal-oriented behavior shaped by information and subject to constraints. However, the actual cognition and behavior of machines differ from those of humans. Machines’ behavior tends to be predictable and persistent 22, with higher precision and faster execution 23, better informed with access to global information 24, and less adaptable and susceptible to influence 25, 26, 27. In contrast, humans tend to be limited to local information, satisfice, act with errors, learn and adapt, succumb to social influence and peer pressure, yet also exhibit opinion stubbornness and behavioral inertia; on occasions, they may also use metacognition and revise their own perceptual and decision-making models. Humans often exhibit cognitive biases due to limited information processing capacity, bounded rationality, reliance on heuristics, vestiges of evolutionary adaptation, and emotional motivations 28, 29, and algorithms trained on data generated by humans may reproduce these biases 30, 31. Research on human-like general AI aims to erase the cognitive and behavioral differences between humans and machines, while work on human-competitive AI strives for superintelligence that is faster, smarter, and more precise than humans’ 32, 33. Either way, humans will remain distinct from machines in the near future.
Research from the CASA (computers as social actors) paradigm in psychology emphasizes that humans treat and respond to machines similarly to other humans: people reciprocate kind acts by computers 34, treat them as politely as they treat humans 35, consider them as competent, but also apply gender and racial stereotypes to them 36, 37; people also humanize and empathize with machines, experiencing distress when witnessing the mistreatment of a robot 38, 39.
Nevertheless, there are visible neurophysiological differences in the brain when humans interact with robots 40, 41, likely because humans do not attribute agency and morals to them 42, 43. AI is perceived to have lower intentional capacity, lack self-interest, and be more unbiased than humans. Consequently, humans exhibit a narrower emotional spectrum with machines than with humans, reacting with lower and flatter levels of social emotions such as gratitude, anger, pride, and a sense of fairness 44, 45, 46, 47, yet judging machines more harshly when they commit mistakes, cause harm, or incur losses 48, 49. Further, humans behave more rationally and selfishly with machines, cooperating and sharing less and demanding and exploiting more 50, 51, 52, 53. People would design a machine to be more cooperative than they are themselves 54 but act pro-socially towards it only if it is more human-like 55, or if it benefits another human 56. Compared to a single person, small groups of people are even more likely to exhibit competitive behavior and bullying toward robots 57. Despite this intergroup bias, humans are still susceptible to machine influence when making decisions or solving problems 58. Robots can cause both informational and normative conformity in people 59, 60 and AI and ChatGPT can corrupt humans’ moral judgment and make them follow unethical advice 61, 62. Humans tend to trust algorithmic advice more than advice coming from a another human or a human crowd 63, 64 but may also avoid it if they perceive a threat to their decision control or a lack of understanding and cognitive compatibility 65, 66.
Collective outcomes
In complex social systems, individuals’ behavior and interactions affect the collective outcomes, but the relationship can be fundamentally different from a simple sum or average 67, 68, 69, 70. The collective outcomes in human-machine social systems differ from those in human-only systems because machines behave differently from humans, human-machine (H-M) and machine-machine (M-M) interactions differ from human-human (H-H) interactions, but also the humans, the machines, and their interactions influence each other indirectly (Fig. 2). We synthesize common dynamics and patterns in groups and networks of humans and machines for five different social interaction situations (Table 1).
Competition
Competition occurs when multiple actors strive for a common goal that cannot be shared, as is the case in contests, auctions, and product markets. Market algorithms are typically designed to benefit the owner without regard for others or the efficiency and stability of the market, and yet, they may still benefit the collective.
With more advanced data processing, learning, and optimization capabilities than humans, algorithms are better able to discover arbitrage opportunities and, hence, eliminate mispricing and increase liquidity in markets. Experimental studies show that algorithmic traders can increase market efficiency 71, but possibly at the expense of human traders’ performance 72, 73. Furthermore, algorithmic traders affect human behavior indirectly: with their presence, they make human traders act more rationally, and thus, reduce strategic uncertainty and confusion in the market 74, reducing price bubbles and bringing prices closer to the fundamental value 75.
While perfectly optimizing arbitrage algorithms eliminate mispricings, neither zero-intelligence algorithms that submit random bids without profit maximization 76 nor profit-maximizing agents that update their beliefs from trading history 77 can improve market quality. Meanwhile, manipulator and spoofing algorithms that act to mislead and influence other traders worsen market efficiency 78. Algorithms can also reduce the rationality of professional traders and alienate and drive away amateur ones. In an online cryptocurrency marketplace, traders herd after a bot buys, producing larger buying volumes 79, while in a Chinese peer-to-peer lending platform, automated investment piques inefficient investor scrutiny, increasing bidding duration without improving investment return 80. In online auction markets, naive first-time bidders respond negatively to being outbid by sniping algorithms and become less likely to return to another auction 81. Snipers, which place last-moment bids 82, work mainly because they exploit the naivety of amateur online bidders, who tend to increase their bids incrementally. However, human lack of rationality has its benefits because squatting (placing a high early bid) deters new entrants 83. In fact, sniping algorithms yield a negligibly small 83 or non-existent 84 buyer gain, giving them a net negative impact on the marketplace.
In addition to trading and auction markets, pricing algorithms have become widespread in regular product markets 85, 86 because they either provide recommendations to human pricing managers 87, 88 or entirely dictate pricing for some firms 86, 89. While pricing algorithms can help firms scale and respond to changes in demand, they may also generate anti-competition. Q-learning algorithms learn to set anti-competitive prices without communication in simulations 90, 91, 92, 93, and in experiments, those algorithms are often more collusive than humans in small markets 94 and foster collusion when interacting with humans compared to fully human markets 95. Observational studies of gasoline markets 89 and e-commerce 96, 97 support the experimental evidence. Furthermore, algorithms can weaken competition by providing better demand predictions, thereby stabilizing cartels 98, 99, 100 or by asymmetries in pricing technologies and commitment 101, 102.
The general intuition is that markets with more actors should be more efficient. Thus, one expects enhanced performance from markets populated by algorithms. However, in reality, the beneficial effects of machines are often in balance, crucially depending on the machines’ prevalence, decision speed, and information quality 103, as well as the humans’ experience and expectations.
Coordination
The problem of coordination requires adopting a strategy identical to or, in some cases, dissimilar from other people’s strategies, as when deciding whether to join a protest, agreeing on a convention such as driving on the right-hand side of the road, adopting a communication technology, or avoiding a crowd or traffic congestion 104, 105.
In human-machine systems, bots could be used to introduce more randomness and movement to steer human groups toward better solutions. Thus, bots acting with small levels of random noise and placed in central locations in a scale-free network decrease the time to coordination, especially when the solutions are hard to find 106. The bots reduce unresolvable conflicts not only in their direct interactions but also in indirect H-H interactions, even when the participants are aware that they are interacting with machines. Bots that are trained on human behavior, however, process information less efficiently and adapt slower, causing hybrid groups playing a cooperative group-formation game to perform worse than human-only and bot-only groups 27.
In sum, in situations where a group may get stuck on a suboptimal equilibrium, non-human-like bots may be able to help by jittering the system with randomness and unpredictability. Such simple bots may be more beneficial than bots that superficially imitate human behavior without the ability to learn and adapt.
Cooperation
The problem of cooperation pertains to social dilemma situations where a decision is collectively beneficial but individually costly and risky. Although the economically rational decision in non-repeated anonymous interactions is to free-ride and exploit others’ contributions, people’s actual behavior tends to be informed by norms of reciprocity, fairness, and honesty signaling. Thus, as a result of millennia of evolutionary adaptation, people generally cooperate with each other. If people know they are interacting with bots, however, they cooperate less 51, 52. Yet, since humans reciprocate to and imitate cooperative neighbors, introducing covert, persistently cooperating bots could increase cooperation.
Computer simulations show that persistent prosocial bots favor the emergence of fairness and cooperation 107, 108, with stronger effects when humans are more prone to imitation and bots occupy more central positions in networks with highly heterogeneous connectivity 109.
Just a few under-cover cooperative bots can increase cooperation, especially if the bots are widespread in the network, interacting with many human players rather than concentrated with overlapping sets of partners 110. The reason is that humans wait for someone else to cooperate before they do, but once they observe many cooperators, they become more likely to exploit.
Yet, cooperative bots may sometimes fail to improve cooperation. For instance, hybrid groups with identifiable bots do not perform better than human-only groups 111. When participants are aware of the presence of artificial agents but not their identity, there is a small increase in cooperation of the bots’ direct neighbors but no significant boost in the overall network 112. Similarly, multiple well-dispersed covert bots, whether all-cooperating or reciprocating, fail to improve cooperation 113, although a single overt network engineer bot who suggests connecting cooperators and excluding defectors can successfully do so.
In sum, covert, persistently cooperating bots (i.e., not very human-like) can increase cooperation in the group depending on the network of interactions. Bots are successful if they are strategically positioned – well dispersed in regular and random networks or centrally located in networks with skewed degree distributions – or have the power to strategically engineer the network by offering opportunities to break links to defectors.
Contagion
Contagion concerns the spread of information and behaviors, such as memes, slang, fashion, emotions, and opinions in communication networks 114, 115, 116. In contrast to the strategic interdependence under the competition, coordination, and cooperation scenarios, the main mechanism here is social influence: the tendency to rely on information from others to handle uncertainty and to conform to the expectations of others to fit in society 117, 118, 119. In human-machine systems, bots can be remarkably influential at the collective level despite exerting limited direct influence on individuals because, in networks, small effects can produce chain reactions and trigger cascades 120, 121, 122, 123.
This is how social media bots influence public opinion. In agent-based models of belief formation, weak bots do not alienate their followers and their followers’ friends and thus have their message spread farther than messages by more pushy and assertive users 124. In other words, network amplification occurs through bots’ indirect influence precisely because their direct influence on humans is weak, slow, and unobtrusive. If social media bots influence not people’s opinion but their confidence to express it, they can amplify marginal voices by triggering the spiral of silence amongst disagreeing humans 26. The bots are more influential when they are more numerous and connected to central actors. Strategically placed zealot bots can in fact bias voting outcomes and win elections 125.
Bots can also trigger emotional contagion in groups, even though they evoke flatter emotional reactions from individual humans. Humanoid robots can encourage and increase social interactions among older adults within care facilities, between different generations, and for children with ASD 57. In small-team collaborative experiments, a robot’s verbal expressions of vulnerability can show “ripple effects” and make the humans more likely to admit mistakes, console team members, and laugh together 126, engage in social conversations and appreciate the group 127. The reported positive contagion effects, however, were detected when comparing one machine to another 127, 128. Overall, bots are more effective than no bots to influence opinions, behavior, and emotions, but not necessarily more effective than humans. Yet, even when bots have a weak direct influence on humans’ opinions, they can exert significant collective influence via persistence, strategic placement, and sheer numbers.
Collective decision-making
Collective decision-making involves groups making choices or solving problems by combining individual opinions. It impacts social phenomena as diverse as team collaboration, voting, scientific innovation, and cultural evolution 129. Originating with Galton’s work on estimation tasks 130, the “wisdom of crowds” concept suggests that a crowd’s aggregated estimate is often more accurate than any individual’s, or sometimes even experts’ 131. Crowds perform better when individual opinions are either independent or diverse 132, while social interaction can hinder 133, 134, 135 or improve 136, 137 collective performance. In human-machine systems, algorithms introduce diversity and can thus improve decision-making.
An analysis of professional Go players’ moves over 71 years suggests that AlphaGo, the AI program Google DeepMind introduced in 2016, led human players to novel strategies and improved their decision-making 138, 139, 140. AlphaGo’s decisions, untethered by human bias, sparked human innovation in this game. However, the positive influence of machine-human social learning on problem-solving may be limited. When algorithms are introduced in chains of humans engaged in sequential problem solving, the innovative solutions benefit immediate followers, but team accuracy does not have lasting effects because humans are more likely to replicate human solutions than algorithmic ones 141. Similarly, in a team prediction task, an algorithm maintaining group diversity by promoting minority opinions improves individual accuracy, but the effects dissipate for team accuracy 142.
The area of hybrid intelligence investigates how and when to combine human and algorithmic decision-making 143, 17 and includes research on active and interactive learning and human-in-the-loop algorithms 144, with applications in clinical decision-making, where combining clinician and algorithmic judgments can improve cancer diagnoses 145, 146, and citizen science, where combining crowd-based with machine classifications can improve accuracy. On Zooniverse, a prominent citizen science platform, this hybrid approach found supernovae candidates among Pan-STARRS images more effectively than humans or machines alone 147 but damaged citizen scientists’ retention 148, 149, suggesting a trade-off between efficiency and volunteer engagement. Ultimately, the deployment of machines could further marginalize certain groups of volunteers 150, and with fewer volunteers, AI’s performance could diminish.
The emerging field of hybrid intelligence suggests that algorithms introduce novel solutions, but these may be too unfamiliar for humans to adopt. Nevertheless, machine diversity and competition might inspire alternative forms of human creativity and innovation. Developing methods to effectively combine human and machine solutions could further improve collective intelligence 151.
Implications
The algorithms in current human-machine social systems are relatively simple. Few use sophisticated machine learning or AI, and typically, these guide narrow and specific behaviors 152, 153. Except for malicious social bots and customer-service chatbots, most machines do not mimic human qualities. Most machines are superhuman — processing vast data amounts, acting swiftly, and handling tedious tasks— or candidly non-human— resisting peer influence, not reciprocating, and acting randomly. There is a clear distinction between covert and overt bots: covert bots are more problematic than bots declaring their identity and following norms and regulations.
The effects of machines on human-machine social systems vary by their number, algorithms, network position, interaction situation, institutional regulations, technological affordances, organizational context, and emerging norms and culture (see Boxes 1-4). Machines alter outcomes through their unique behavior because humans interact differently with them, and because of their indirect effects – machines’ presence changes how humans interact amongst themselves.
Machines can be beneficial when they act or steer humans to counteract human weaknesses. For instance, noisy bots can disrupt sub-optimal outcomes and improve coordination, persistently cooperative bots can curb retaliation and maintain cooperation, machines in central roles as arbitrageurs improve price discovery and market quality, and network-engineering bots boost collective welfare via cooperator assortment and defector exclusion. With global information, higher processing power, and instantaneous execution, machines can quickly address external events like vandalism or political and natural crises, ensuring system robustness, resilience, and efficiency. Depending on the situation, machines offer superhuman persistence or randomness, norm-setting rationality, or solution diversity, enhancing human behavior towards better outcomes.
However, what helps in one context can hinder in another. Machines’ unintuitive solutions may confuse humans, hindering innovation and technological progress. Humans might not act fast enough to correct machines’ errors, resulting in instabilities and flash failures. Machines are less adaptive than humans to changes, impeding system evolution. Machines can be designed to exploit human weaknesses, triggering cascades that exacerbate polarization, emotional contagion, ideological segregation, and conflict. Machines’ non-human optimization logic, execution speed, and behavioral rigidity can clash with human behavior, pushing interactions toward undesirable outcomes.
Research
Existing research is often biased towards engineering and optimization, lacking deeper insights from a social science perspective. The time for a new sociology of humans and machines is critical, before AI becomes more sophisticated: generative AI exhibits emergent behavior that itself requires explanation 154, 155, complicating the understanding of system dynamics.
Researchers would benefit from an agent-based modeling framework that outlines distinctions between human and bot agents: utility function, optimization ability, access to information, learning, innovation/creativity, accuracy, etc. The framework could borrow concepts from other two-agent systems, such as predator–prey, principal–agent, and common pool resource models. Controlled experiments should explicitly compare human-machine, human-only and machine-only networks, and known bots against covert bots. Experiments could manipulate participants’ perceptions of algorithms’ technical specifications, agenthood 156, emotional capability, and biases. Field interventions in online communities with endemic bot populations present another promising direction. Existing examples include social bots that gain influence by engaging human users 157, 158, 159, 160, trading bots that manipulate prices in cryptocurrency markets 161, political bots that promote opposing political views to decrease polarization 162, and “drifters” to measure platform bias 80. Expanding on the cases reported here, we need observational research on additional human-machine communities and contexts such as traffic systems with human-driven and driverless vehicles, online multiplayer games comprising human players, non-player characters, and cheating code, and dating markets with AI-driven chatbots 163.
Finally, research with artificial agents introduces ethical problems demanding careful elaboration and mitigation. Research protocols should minimize interventions 164, possibly deploying covert bots only where they already exist, ensuring their actions are not unusual or harmful 79. Even then, bots may still face opposition from users due to privacy concerns 157. Do people perceive certain bots as inherently deceptive? Could knowledge of the bot owner and algorithm mitigate this perception?
Design
Humans are resilient and successful when they exhibit high levels of efficient communication, context awareness, emotional recognition and response, and ethical and cultural sensitivities. These features should be encouraged when designing social machines to build trust, ensure legal compliance, promote social harmony, enhance user satisfaction, and achieve long-term sustainability 165, 166.
Since H-H, H-M, and M-M interactions differ, machines should be specifically designed for each scenario, with separate training for each interaction. This could avoid market underperformance when bargaining algorithms trained on human-only markets adapt poorly to human-machine negotiations 50, or traffic jams when driverless vehicles trained on human driving fail to properly interact with each other 167.
AI design should also adopt a hierarchy of behavioral rules and conventions guiding H-M and M-M interactions in the context of H-H interactions. Isaac Asimov’s famous Three Laws of Robotics 168 which regulate M-H interactions and self-preservation—1) not harming humans, 2) obeying humans, and 3) protecting own existence, with priority given to higher order rules—could be adapted for M-M interactions, considering specific contexts, implications, and unintended consequences 169.
Further, cultural context is crucial for AI design. People’s perceptions of machines vary by age, environment, personality, and geography 170, 46. Machines reflect their developers’ culture and operate in settings with specific organizational and community norms 22. Thus, AI design for self-driving cars and assistant bots should consider local driving culture and attitudes toward domestic assistants during training.
Finally, the systems we reviewed exhibit complexity in that their emergent behavior transcends a mere aggregation of individual components, displaying what is known as “network effects.” While natural complex systems often demonstrate remarkable resilience and adaptivity 171, the human-designed complex systems discussed here, albeit to varying degrees, are not inherently adaptive. To enhance resilience and robustness, AI designers should incorporate complex adaptive system principles, like negative feedback, modularity, and hierarchical organization. For example, dense networks lacking diversity and modularity are susceptible to systemic failures 172, 173, 174 and are easy to control 175. Introducing bots that increase network diversity, introduce resistance, build resilience, or incorporate negative feedback could enhance adaptability and stability. Such configurations have improved group outcomes in forecasting 176, exploration/exploitation 177, and general knowledge tasks 137. Literature on monitoring and steering complex systems can help predict critical transitions 178, 179, 15, designing safer human-machine systems across different environments and perturbations.
Policy
We urge a system-focused approach to AI policy and ethics: policymakers should approach AI not as a single existential threat but as a multiplicity of machines and algorithms. Machines are often more beneficial when they are superhuman or simply “alien” and when they are diverse. Similarity in information sources, interaction speeds, optimization algorithms, and objective functions can cause catastrophic events, like flash crashes in markets. Thus, while AI designers may chase optimization and superintelligence, policymakers should focus on the diversity of human-machine ecologies. Policymakers should demand adaptivity and resilience, too.
Policymakers should also anticipate the social co-development of machines and humans, which will inadvertently change existing institutions. Machines can cause humans to withdraw interaction: for instance, outsourcing care to robots reduces caregivers’ empathy 180. Intelligent machines are changing the transmission and creation of human culture, altering social learning dynamics, and generating new game strategies, scientific discoveries, and art forms 181. Humans must adapt to intelligent machines just as intelligent machines must learn from and adapt to humans. Finally, ethicists should address questions such as: Should all machines be equal? Should we allow status hierarchies, possibly reflecting and exacerbating existing socio-economic inequalities?
Conclusion
This Perspective synthesizes relatively disparate literature based on agent-based models, controlled experiments, online field interventions, and observational analyses from human-computer interaction, robotics, web science, financial economics, and computational social science under a common theoretical framework: human-machine social systems. We identify common dynamics and patterns that emerge from the interactions of humans and intelligent machines regardless of the specific context, as well as peculiarities and unique problems that concrete techno-organizational and socio-cultural environments generate. Our utmost ambition is to stimulate cumulative empirically driven and mechanism-focused sociological research in the emerging, fast-evolving field of human-AI science. At stake are new and urgent social challenges such as online misinformation, market flash crashes, cybersecurity, labor market resilience, and road safety. With increasing social connectivity and accelerating developments in AI, understanding the complex interactions between humans and intelligent machines is a challenging undertaking, but one that is crucially important for a better human future.
Acknowledgments
MT acknowledges the generous support of the Santa Fe Institute during the period when the research was conducted. TY was partially funded by the Irish Research Council under grant number IRCLA/2022/3217, ANNETTE (Artificial Intelligence Enhanced Collective Intelligence). The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.